# System-Theoretic Analysis of Dynamic Generalized Nash Equilibrium Problems – Turnpikes and Dissipativity

### v1

@@ 3.1 Dissipativity of GNEPs
$\text{s.t.}\;\bar{x}=f(\bar{x},\bar{u}),\;g(\bar{x},\bar{u})\leq 0,\;G\,\bar{u}\leq h.$
Substituting it into (11) yields
$\ell(x_{s},u_{s})+\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{\diamond}-x_{s}\\ u^{\diamond}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x^{\diamond},u^{\diamond}),$
where $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$. This inequality, however, holds only if $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$ which implies that the steady-state GNE (10) has to coincide with the minimizer of (12), and thus be a solution to an optimization problem. In turn, this would require the game to be potential [Monderer and Shapley, 1996] and the potential function to be the sum of agents’ costs [Slade, 1994, Prp. 1]. More specifically, we see from the first-order optimality conditions that it would require
$\frac{\ell^{v}(x_{k},u_{k})}{\partial x_{k}}=\frac{\ell(x_{k},u_{k})}{\partial x_{k}}$ (13a)
$\frac{\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}=\frac{\ell(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}$ (13b)
and as $\ell(x_{k},u_{k})=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k})$ this would impose that all agents’ cost functions are the same up to a constant which does not hold for the GNEPs we consider in (2). It is a well known phenomenon that selfish agent behavior in games leads to solutions that are suboptimal compared to centrally minimizing the sum of costs (Price of Anarchy). [Kulkarni, 2019] analyze these efficiency losses for GNEs. Thus, in Definition 3 we require the dissipation inequality to hold only for trajectories resulting from the GNEP in (2).
The next result shows that strict dissipativity also provides an avenue to analyze which $(x_{S},u_{s})\in\mathcal{S}^{GNE}_{S}$ can appear in (11).
Lemma 2.
Let the strict dissipation inequality (11) hold for all $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$. Then
$\sum_{v\in\mathcal{V}}\ell^{v}(x_{s},u_{s}^{v})=:\ell(x_{s},u_{s})<\ell(\bar{x},\bar{u}):=\sum_{v\in\mathcal{V}}\ell^{(}\bar{x},\bar{u}^{v})$
holds for all $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$.
Proof 3.1.
For contradiction, suppose there exists a better steady state solution $(\bar{x},\bar{u})\in\mathcal{S}_{s}^{GNE}$ with $\ell(x_{s},u_{s})>\ell(\bar{x},\bar{u})$. At steady state, the strict inequality (11) gives
$0<\alpha_{\ell}\left(\left\|\begin{matrix}\bar{x}-x_{s}\\ \bar{u}-u_{s}\end{matrix}\right\|\right)\leq\ell(\bar{x},\bar{u})-\ell(x_{s},u_{s}).$
Since $\alpha_{\ell}\in\mathcal{K}$, this allows writing $\ell(x_{s},u_{s})<\ell(\bar{x},\bar{u})$ and thus shows the contradiction.
The last result requires (11) to hold on (10). This is needed, as without further analysis it is not clear that for $x_{0}=x_{s}$ the GNE control action at time $k=0$ is such that $u^{*}_{0}=u_{s}$. Indeed, the control action $u_{0}^{*}$ depends on the specific design of (2) and the specific horizon $N$. Corollary 10 in Section 4 presents a result that ensures $u^{*}_{0}=u_{s}$ if $x_{0}=x_{s}$.
We now formally introduce the considered turnpike property for our GNE problem in (2) with respect to the steady-state GNE.
@@
The GNEP (2) exhibits the (measure) turnpike property at $(x_{s},u_{s})$ if for each $\varepsilon>0$ there exists $C>0$ such that $\forall N\in\mathbb{N},\mathbf{x}\in\mathbb{X}_{0}$ and for all game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ it holds that
$Q_{\varepsilon}:=\#\{k\in\mathbb{Z}_{N}|\,\|x_{k}-x_{s}\|\leq\varepsilon\}\geq N-\frac{C}{\alpha_{\ell}(\varepsilon)}.$ (14)
We next show that strict dissipativity of the GNEP with respect to $(x_{s},u_{s})$ implies the turnpike property for the input and state trajectories resulting from (2). The result is inspired by [Grüne, 2013, Thm. 5.3] and its proof.
Theorem 3 (Str. dissipativity $\Rightarrow$ turnpike).
If, for every $x\in\mathbb{X}_{0}$, the GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ in the sense of Definition 3, then all GNEP solutions with $J_{N}(x,u)\leq N\ell(x_{s},u_{s})+\delta$ exhibit the turnpike property at $(x_{s},u_{s})$.
Figure 2: Overview of implications between strict dissipativity, turnpike and optimal operation at the steady-state GNE.
Proof 3.2.
We define $C:=2\sup_{\mathbf{x}\in\mathbb{X}_{0}}|\Lambda(\mathbf{x})|<\infty$ and further we have that
$\tilde{J}_{N}(x,u)$ $=J_{N}(x,u)+\Lambda(\mathbf{x})-\Lambda(x_{N})$ (15a)
$\leq J_{N}(x,u)+C,$ (15b)
$\leq N\,\ell(x_{s},u_{s})+(\delta+C)$ (15c)
@@
$\ell(x_{k},u_{k})+\Lambda(x_{k})-\Lambda(f(x_{k},u_{k}))\geq$
$\ell(x_{s},u_{s})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)\$(16)
For contradiction suppose that the trajectory does not fulfill the turnpike property and thus, $Q_{\varepsilon}<N-\frac{\delta+C}{\alpha_{\ell}{(\varepsilon})}$. This implies that there exists a number of $N-Q_{\varepsilon}>\frac{\delta+C}{\alpha_{\ell}(\varepsilon)}$ times steps in the set $\mathcal{N}\subseteq\{0,\dots,N-1\}$ for which the trajectory is more than $\varepsilon$ distance away from the steady state $x_{s}$, i.e., $\|x_{k}-x_{s}\|>\varepsilon,\;\forall k\in\mathcal{N}$.
$\tilde{J}_{N}(x,u)$ $=\sum_{k=0}^{N-1}[\ell(x_{k},u_{k}))+\Lambda(x_{k})-\Lambda(f(x_{k},u_{k}))]$
@@
$=N\ell(x_{s},u_{s})+Q_{\varepsilon}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)]$
$+\underbrace{(N-Q_{\varepsilon})\,\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)}_{\geq(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)}$
@@
Theorem 3 demonstrates that strict dissipativity of the GNEP (2) is not merely a theoretical property but gives direct insights into the behavior of the trajectory with respect to the steady-state GNE.
@@
Observe that—similar to its optimal control counterpart—the proof of Theorem 3 does not rely on $(x,u)$-dissipativity but only requires $x$-dissipativity with $\alpha_{\ell}(\|x-x_{s}\|)$. However, having strictness in $x$ and $u$ helps to derive converse turnpike results.
@@ 3.2 A converse turnpike result for GNEPs
The question arises if the converse also holds, i.e., whether the turnpike behavior in the open-loop trajectories implies strict dissipativity with respect to $(x_{s},u_{s})$. We recall the following “off-equilibrium” concept.
Definition 5 ($\varepsilon$-GNE [Chen et al., 2021]).
@@
For a given $\varepsilon\geq 0$, $u^{*}$ is said to be an $\varepsilon$-generalized Nash equilibrium ($\varepsilon$-GNE) of (2) if $\forall v\in\mathcal{V}$ and $\forall(x,u^{v})\in\mathcal{Z}(\mathbf{x},u^{-v*})$ it holds:
$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})+\varepsilon,\;$
@@
The solution set is denoted as $\mathcal{S}^{\varepsilon-\text{\tiny GNE}}_{N}(\mathbf{x})$. Particularly, when $\varepsilon=0$, $u^{*}$ is a GNE and $\mathcal{S}^{\varepsilon-\text{\tiny GNE}}_{N}(\mathbf{x})=\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
@@
We make the following assumption which ensures that $(x_{s},u_{s})$ is the unique GNE in the set of $\rho$-GNEs centred at $x_{s}$ , $\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})$, which minimizes the performance measure $J(x,u)$ defined for the population of agents.
Assumption 1 (Local cost minimizer $(x_{s},u_{s})$).
@@
There exists a constant $\rho$ and $\alpha_{\rho}\in\mathcal{K}$ such that $\forall(x,u)\in\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})\cap\mathcal{Z}(x_{s})$ and $\forall k\in\mathbb{N}$ it holds:
$\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x_{k},u_{k})-\ell(x_{s},u_{s}).$ (17)
@@
Assumption 1 can be verified by employing the game-theoretic counterpart of Fiacco-like sensitivity analysis. Specifically, in the GNE setting formulated as quasi-variational inequalities, local uniqueness results exist under different sets of assumptions, see [Alphonse et al., 2021, Ch. 3] and [Noor et al., 2025, Lem. 7]. For variational GNEs, sensitivity analysis of variational inequalities can be exploited [Facchinei and Pang, 2009, Prop. 12.14]. The following result confirms a converse relation between turnpike and strict dissipativity.
Theorem 4 (Turnpike $\Rightarrow$ str. dissipativity).
@@
Suppose at $(x_{s},u_{s})$ Assumption 1 and the turnpike property as per Definition 4 holds for the GNEP (2), then the GNEP is strictly dissipative for the supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ with respect to $(x_{s},u_{s})$.
Proof 3.3.
Consider $\alpha_{\rho}\in\mathcal{K}$ from Assumption 1 and note that for any $\rho$ there exist an $N\in\mathbb{N}$ such that $Q_{\varepsilon}>0$ as defined in (14). Consequently, we choose $\rho\geq\varepsilon$ and thus whenever we are in $Q_{\varepsilon}$ we can exploit Assumption 1.
Recall the available storage from Theorem 1
$\tilde{\Lambda}_{a}(\mathbf{x})=\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$ (18)
@@
evaluated along the game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$. Temporarily, consider some $N\in\mathbb{N}$ and a feasible initial condition $\mathbf{x}$. We split the time horizon into $Q_{\varepsilon}(N,\mathbf{x})$, the trajectory points close to $(x_{s},u_{s})$ and $Q_{out}(N,\mathbf{x})=\{0,\dots,N-1\}\backslash Q_{\varepsilon}(N,\mathbf{x})$ points outside of ball of radius $\varepsilon$. Evaluating (18) on $Q_{out}(N,\mathbf{x})$ gives
@@
where $\hat{\alpha}=\sup_{(x,u)\in\mathcal{Z}(\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ and similarly $\hat{s}=\sup_{(x,u)\in\mathcal{Z}(\mathbf{x})}s(x_{k},u_{k})$. The Turnpike property implies that any trajectory, independent of $N$ and $\mathbf{x}$, only spends a finite amount of time outside of ball the of radius $\varepsilon$ of $x_{s}$ which allows us to establish the following bound:
@@
Next, we evaluate the second part of the trajectory for $Q_{\varepsilon}(N,\mathbf{x})$ we have that $\|x_{k}-x_{s}\|\leq\rho,\;\forall k\in Q_{\varepsilon}(N,\mathbf{x})$ and given Assumption 1 it holds that:
$\sum_{Q_{\varepsilon}(N,\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})\leq 0$ (21)
Combining (3.3) and (21) we get that

### v3

@@ 3 Dissipativity and Turnpikes for GNEPs
$\text{s.t.}\;\bar{x}=f(\bar{x},\bar{u}),\;g(\bar{x},\bar{u})\leq 0,\;h(\bar{u})\leq 0.$
in which $\ell(\bar{x},\bar{u})=\displaystyle\sum_{v\in\mathcal{V}}\ell^{v}(\bar{x},\bar{u}^{v},u^{-v})$ as defined in (12) and $h(\cdot)$ stacks all local constraints. Consider the dissipativity inequality in (14)
$\Lambda(f(x_{k},u_{k}))-\Lambda(x_{k})\leq s(x_{k},u_{k})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$
which evaluated at $(x^{\diamond},u^{\diamond})$ for $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ as given in Definition 3 yields
$\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{\diamond}-x_{s}\\ u^{\diamond}-u_{s}\end{smallmatrix}\right\|\right)\leq\ell(x^{\diamond},u^{\diamond})-\ell(x_{s},u_{s})$ (17)
where $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$ is the steady-state GNE. By construction, $(x^{\diamond},u^{\diamond})$ minimizes $\ell(\bar{x},\bar{u})$. Hence (17) holds only if $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$, i.e., the steady-state GNE must coincide with the minimizer of (16). This, however, does not hold in general for the GNEP in (2) which asserts our claim.
Example 1 (When do GNEP and OCP coincide?).
We now provide intuition for when the price of anarchy (14) equals one, focusing on the steady-state problem (11) for simplicity. Indeed, one can identify conditions when $(x_{s},u_{s})=(x^{\diamond},u^{\diamond})$ by comparing the first-order optimality conditions of (11) and (16); they coincide if
$\frac{\partial\ell^{v}(x_{k},u_{k})}{\partial x_{k}}=\frac{\partial\ell(x_{k},u_{k})}{\partial x_{k}}$ (18a)
$\frac{\partial\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}=\frac{\partial\ell(x_{k},u_{k}^{v},u_{k}^{-v})}{\partial u^{v}_{k}}$ (18b)
holds $\forall v\in\mathcal{V}$ and as $\ell(x_{k},u_{k})=\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k})$ this would require for all agents’ coupled cost function parts to be the same up to a constant. This would correspond to a potential game [46] with the sum of costs as the potential function [53, Prp. 1]. We do not impose this restriction in (2).
We now formally introduce the considered turnpike property for our GNE problem in (2) with respect to the steady-state GNE (11).
@@
The GNEP (2) exhibits the (measure) turnpike property at $(x_{s},u_{s})$ if for each $\varepsilon>0$ there exists $C>0$ such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$, and for all game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ it holds that
$Q_{\varepsilon}:=\#\left\{k\in\mathbb{Z}_{N}|\,\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\leq\varepsilon\right\}\geq N-\frac{C}{\alpha(\varepsilon)}$ (19)
for some $\alpha\in\mathcal{K}$ and where #\# refers to the cardinality.
Definition 4 provides a lower bound on the number of time steps the GNE trajectory spends within $\mathcal{B}_{\varepsilon}(x_{s})$, an $\varepsilon$-ball around the steady-state GNE. This bound grows with the horizon $N$, thus intuitively as $N$ increases, the trajectory spends more time in a neighborhood around $x_{s}$, while the time spent outside $\mathcal{B}_{\varepsilon}(x_{s})$ remains bounded independently of $N$.
Assumption 1 (Cheap reachability).
For all $\mathbf{x}\in\mathbb{X}_{0}$, there exists an infinite-horizon feasible pair $(x,u)\in\mathcal{Z}_{\infty}(\mathbf{x})$ such that, for some $\delta\in\mathbb{R}$ and all $N\in\mathbb{N}$ it holds that
$J_{N}(x,u)\leq N\ell(x_{s},u_{s})+\delta.$
It is worth noting that this assumption is readily verified if from all $\mathbf{x}\in\mathbb{X}_{0}$, the equilibrium point $x_{s}$ can be reached in a finite number of steps or if the Jacobian linearization of (1) at $(x_{s},u_{s})$ is stabilizable. Moreover, we note that the feasible $(x,u)$ in Assumption 1 does not need to be a game pair.
Assumption 2 (Bounded price of anarchy).
For all $\mathbf{x}\in\mathbb{X}_{0}$, the price of anarchy from (14) satisfies
$\text{PoA}(\mathbf{x})=\frac{\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})}{V^{\diamond}_{N}(\mathbf{x})}\leq P<\infty$
for $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and $0<\nu\leq V^{\diamond}_{N}(\mathbf{x})\leq V<\infty$ holds.
We next show that strict dissipativity of the GNEP with respect to $(x_{s},u_{s})$ implies the turnpike property for the input and state trajectories resulting from (2). The result is inspired by [34, Thm. 5.3] and its proof.
Theorem 3 (Strict dissipativity $\Rightarrow$ turnpike).
Consider the GNEP (2) and let Assumptions 1 and 2 hold. Suppose that the GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ in the sense of Definition 3 and the storage is bounded on $\mathbb{X}_{N}(\mathbb{X}_{0})$. Then all GNEP solutions exhibit the (measure) turnpike property at $(x_{s},u_{s})$.
Proof 3.3.
The proof proceeds in two steps. First we show that Assumptions 1 and 2 allows to bound the performance of the population of agents along game pairs. Second we exploit the strict dissipation inequality to construct the asserted bound (19).
Step 1:  Lemma 39 from Appendix B shows that Assumption 2 implies the linear performance bound $J_{N}(x^{*},u^{*})\leq V^{\diamond}_{N}(\mathbf{x})+\tilde{P}$ for some finite $\tilde{P}$. Since Assumption 1 relies on non-optimal trajectories, we have the bound
$J_{N}(x^{*},u^{*})\leq N\ell(x_{s},u_{s})+\delta+\tilde{P}$
which holds for all $N\in\mathbb{N},\mathbf{x}\in\mathbb{X}_{0}$ and $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
Step 2:  Let $C:=2\sup_{\mathbf{x}\in\mathbb{X}_{0}}|\Lambda(\mathbf{x})|<\infty$. Then we have
$\tilde{J}_{N}(x^{*},u^{*})$ $:=J_{N}(x^{*},u^{*})+\Lambda(\mathbf{x})-\Lambda(x_{N})$ (20a)
$\leq J_{N}(x^{*},u^{*})+C,$ (20b)
$\leq N\,\ell(x_{s},u_{s})+(\delta+\tilde{P}+C)$ (20c)
@@
$\ell(x^{*}_{k},u^{*}_{k})+\Lambda(x^{*}_{k})-\Lambda(f(x^{*}_{k},u^{*}_{k}))\geq$
$\ell(x_{s},u_{s})-\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{*}_{k}-x_{s}\\ u^{*}_{k}-u_{s}\end{smallmatrix}\right\|\right)\$(21)
Next we construct the bound from (19) with $\alpha=\alpha_{\ell}$. For contradiction suppose that the trajectory does not fulfill the turnpike property and thus, $Q_{\varepsilon}<N-\frac{\delta+\tilde{C}}{\alpha_{\ell}{(\varepsilon})}$ with $\tilde{C}=C+\tilde{P}$. This implies that there exists a number of $N-Q_{\varepsilon}>\frac{\delta+\tilde{C}}{\alpha_{\ell}(\varepsilon)}$ times steps in the set $\mathcal{N}\subseteq\{0,\dots,N-1\}$ for which the trajectory is more than $\varepsilon$ distance away from the steady state $(x_{s},u_{s})$, i.e., $\|(x^{*}_{k},u^{*}_{k})-(x_{s},u_{s})\|>\varepsilon,\;\forall k\in\{0,\dots,N-1\}$. We obtain
$\tilde{J}_{N}(x^{*},u^{*})$ $=\sum_{k=0}^{N-1}[\ell(x^{*}_{k},u^{*}_{k}))+\Lambda(x^{*}_{k})-\Lambda(f(x^{*}_{k},u^{*}_{k}))]$
@@
$=N\ell(x_{s},u_{s})+Q_{\varepsilon}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{*}_{k}-x_{s}\\ u^{*}_{k}-u_{s}\end{smallmatrix}\right\|\right)]$
$+\underbrace{(N-Q_{\varepsilon})\,\alpha_{\ell}\left(\left\|\begin{smallmatrix}x^{*}_{k}-x_{s}\\ u^{*}_{k}-u_{s}\end{smallmatrix}\right\|\right)}_{\geq(N-Q_{\varepsilon})\alpha_{\ell}(\varepsilon)}$
@@
The proof relies on Assumptions 1 and 2 together with strict dissipativity with bounded storage. Assumption 2, specific to the GNEP setting, requires that the overall GNEP performance is not arbitrarily worse than its OCP counterpart, which is arguably a mild condition. Assumption 1 is not specific to GNEPs as it is a system property. It can be verified via reachability arguments as done in the OCP turnpike literature [34]. Similarly, boundedness of the storage function $\Lambda$ is standard in dissipativity-based turnpike analysis. Actually Theorem 1 shows that the available storage is always bounded from above and below, so assuming strict dissipativity with bounded storage is no stronger than assuming (14) alone. Finally, one might argue that the dissipativity assumption in itself is artificial. Yet, Theorem 3 provides an explanation for the turnpike phenomenon in GNEPs and it elegantly links the dynamic game trajectories to the steady-state GNE. Moreover, we will see in Theorem 4 that also converse statements can be derived: turnpike $\Rightarrow$ dissipativity.
@@
Observe that—similar to its optimal control counterpart—the proof of Theorem 3 does not rely on $(x,u)$-dissipativity with $\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ but only requires $x$-dissipativity with $\alpha_{\ell}(\|x_{k}-x_{s}\|)$. That is, the result of Theorem 3 can be slightly generalized to show a measure turnpike at $x_{s}$ only. However, having strictness in $x$ and $u$ will later help to derive converse turnpike results.
@@ 3.1 A converse turnpike result for GNEPs
We recall the following “off-equilibrium” concept.
Definition 5 ($\rho$-GNE [15]).
@@
For a given $\rho\geq 0$, the pair $(u^{*},x^{*})$ is said to be an $\rho$-generalized Nash equilibrium ($\rho$-GNE) of (2) if $\forall v\in\mathcal{V}$ and $\forall(x,u^{v})\in\mathcal{Z}_{N}(\mathbf{x},u^{-v*})$ it holds:
$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})+\rho,\;$
@@
The solution set is denoted as $\mathcal{S}^{\rho\text{-\tiny GNE}}_{N}(\mathbf{x})$. Particularly, when $\rho=0$, $(u^{*},x^{*})$ is a GNE and $\mathcal{S}^{\rho\text{-\tiny GNE}}_{N}(\mathbf{x})=\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
@@
The following assumption ensures that, among all $\rho$-GNEs centred at $x_{s}$, the pair $(x_{s},u_{s})$ is the one that minimizes the performance measure $J(x,u)$ defined for the population of agents.
Assumption 3 (Local equilibrium cost bound).
@@
There exists a constant $\rho>0$ and $\alpha_{\rho}\in\mathcal{K}$ such that $\forall(x,u)\in\mathcal{S}^{\rho\text{\tiny-GNE}}_{N}(x_{s})\cap\mathcal{Z}_{N}(x_{s})$ and $\forall k\in\mathbb{N}$ it holds:
$\ell(x_{s},u_{s})\leq\ell(x_{k},u_{k})-\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$ (22)
where $\mathcal{Z}_{N}(x_{s})$ is the compact set of all feasible (game) trajectories of length $N$ initialized at $x_{s}$ as in (4).
@@
In principle, one could also consider $\mathcal{S}^{\rho_{1}\text{-\tiny GNE}}_{N}(\mathbf{x})$ with a smaller “suboptimality” gap $\rho_{1}\leq\rho$ which still ensures that Assumption 3 holds for all trajectories in $\mathcal{S}^{\rho_{1}\text{-\tiny GNE}}_{N}(\mathbf{x})$. Assumption 3 can be verified by employing the game-theoretic counterpart of Fiacco-like sensitivity analysis. Specifically, in the GNE setting formulated as quasi-variational inequalities, local uniqueness results exist under different sets of assumptions, see [1, Ch. 3] and [49, Lem. 7]. For variational GNEs, sensitivity of variational inequalities can be exploited [25, Prop. 12.14].
The following result confirms a converse relation between turnpike and strict dissipativity.
Theorem 4 (Turnpike $\Rightarrow$ strict dissipativity).
@@
Suppose that the stage cost $\ell$ is continuous, that at $(x_{s},u_{s})$ Assumption 3, and the turnpike property as per Definition 4 holds for the GNEP (2). Then, there exists some $\alpha\in\mathcal{K}$, such that the GNEP is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ with respect to $(x_{s},u_{s})$. Moreover, there exists a storage function $\Lambda$ bounded on $\mathbb{X}_{\infty}(\mathbb{X}_{0})$.
Proof 3.4.
Consider $\rho$ for which Assumption 3 holds and the corresponding $\alpha_{\rho}\in\mathcal{K}$. The turnpike property in (19) states that given $\rho$ [^1] To improve readability of the proof, we choose $\varepsilon=\rho$ but one may also consider a smaller region around $x_{s}$ (i.e., $\varepsilon\leq\rho$). there exists a horizon length $N\in\mathbb{N}$ for which the GNE trajectory spends nonzero discrete time steps $Q_{\rho}>0$ within $\mathcal{B}_{\varepsilon}(x_{s})$. Further, recall the definition of the available storage from Theorem 1
${\Lambda}_{\alpha_{\rho}}(\mathbf{x})=\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$ (23)
evaluated along the game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$. Here, without loss of generality, we use the $\mathcal{K}$-class function $\alpha_{\rho}$ from Assumption 3.
@@
Temporarily, consider some $N\in\mathbb{N}$ and a feasible initial condition $\mathbf{x}$. We split the time horizon into $Q_{\rho}(N,\mathbf{x})$, the trajectory points close to $(x_{s},u_{s})$, and
$Q_{out}(N,\mathbf{x}):=\{0,\dots,N-1\}\backslash Q_{\rho}(N,\mathbf{x}),$
i.e. the points outside of $\mathcal{B}_{\rho}(x_{s})$. Evaluating (23) on $Q_{out}(N,\mathbf{x})$ gives
@@
where
$\hat{\alpha}=\sup_{(x,u)\in\mathcal{Z}_{N}(\mathbf{x})}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}-x_{s}\\ u_{k}-u_{s}\end{smallmatrix}\right\|\right)$
and similarly $\hat{s}=\sup_{(x,u)\in\mathcal{Z}_{N}(\mathbf{x})}s(x_{k},u_{k})$. The turnpike property implies that any trajectory, independent of $N$ and $\mathbf{x}$, only spends a finite amount of time outside of the ball $\mathcal{B}_{\rho}(x_{s})$. This allows us to establish the bound
@@
where due to (19) $\bar{Q}_{out}\geq\frac{C}{\alpha_{\rho}(\rho)}$ is some constant independent of $N$ and $\mathbf{x}$.
Next, we evaluate the second part of the trajectory. For $Q_{\rho}(N,\mathbf{x})$ we have that due to the turnpike property $\forall k\in Q_{\rho}(N,\mathbf{x})$ it holds that $\|x_{k}-x_{s}\|\leq\rho$. Assumption 3 holds $\forall k\in Q_{\rho}(N,\mathbf{x})$ by choice of $\rho$. Thus we can construct the following inequality:
$\sum_{k\in Q_{\rho}(N,\mathbf{x})}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})\leq 0$ (26)
where we use the fact that $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ . Combining (3.4) and (26) we get that
$\sum_{k=0}^{N-1}\alpha_{\rho}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*})$